OpenStudy (anonymous):

find an equation of the form y-k = a(x-h)^2 for the parabola with vertex (-1,5) and y intercept 8

OpenStudy (anonymous):

Recall: \[y=a(x-h)^{2}+k\] is vertex form for a parabola, where (h,k) is the vertex. So if you need the vertex to be (-1,5), you can plug those values in for h and k. \[y=a(x-h)^{2}+k\] \[y=a(x-(-1))^{2}+(5)\] and simplify \[y=a(x+1)^{2}+5\] From here, we're going to have to manipulate the a value in order to get this thing to have a y-intercept of 8. As it sits right now, the y-intercept is made by taking: \[a(h)^{2}+k\] If you want to design it so that it goes through the y axis at 8, you have to set it equal to 8 and solve for a. \[a(-1)^{2}+5=8\] simplify a+5=8 subtract 5 on both sides a = 3 Now, we have one final manipulation to get it back to the form they asked for it in. \[y=3(x+1)^{2}+5\] You want to put the 5 on the same side as the y, so you can subtract it off of both sides. So, finally: \[y-5=3(x+1)^{2}\] I hope that helps!! :)

OpenStudy (anonymous):

yesss thankss

OpenStudy (anonymous):

If the idea of \[a(h)^{2}+k \] is a little too weird, you can always force the a value out by starting with the equation as we have it at that point: \[y=a(x+1)^{2}+5\] And plugging in the point that we need to be the vertical intercept (0,8) for x and y. \[(8)=a((0)+1)^{2}+5\] If you solve for a here, you'll get the same result. This is a little more traditional too. :) So first, simplify \[8=a(1)^{2}+5\] \[8=1a+5\] \[8=a+5\] Then subtract 5 from both sides, to get: 3=a